# Ye Bai1, Minghan Wang1, Thuy-Trang Vu1

### v1

@@ Abstract
Table-based question answering requires complex reasoning capabilities that current LLMs struggle to achieve with single-pass inference. Existing approaches, such as Chain-of-Thought reasoning and question decomposition, lack error detection mechanisms and discard problem-solving experiences, contrasting sharply with how humans tackle such problems. In this paper, we propose MAPLE  (Multi-agent Adaptive Planning with Long-term mEmory), a novel framework that mimics human problem-solving through specialized cognitive agents working in a feedback-driven loop. MAPLE  integrates 4 key components: (1) a Solver using the ReAct paradigm for reasoning, (2) a Checker for answer verification, (3) a Reflector for error diagnosis and strategy correction, and (4) an Archiver managing long-term memory for experience reuse and evolution. Experiments on WiKiTQ and TabFact demonstrate significant improvements over existing methods, achieving state-of-the-art performance across multiple LLM backbones. [^1] Code and data will be released upon acceptance.
@@ 1 Introduction
Tables represent one of the most prevalent forms of semi-structured data, organizing information systematically across domains ranging from scientific research to business analytics (Dong and Wang, 2024). However, answering questions over tables presents unique challenges, requiring multi-step reasoning over structured data, recognition of implicit relationships between cells, and precise contextual interpretation (Lu et al., 2025). These challenges make table-based question answering (QA) particularly difficult for Large Language Models (LLMs), as they must navigate tabular data structure while performing sophisticated reasoning to derive accurate answers, capabilities that current LLMs struggle to achieve with single-pass inference.
Existing table reasoning frameworks exhibit several limitations. Single-forward-pass methods (Cheng et al., 2023; Ye et al., 2023) lack error detection mechanisms, allowing mistakes to propagate through solutions. ReAct-based approaches (Wang et al., 2024b; Zhang et al., 2023) provide environmental feedback but lack systematic verification. On the other hand, multi-agent approaches primarily focus on output refinement rather than comprehensive reasoning improvement (Ye et al., 2023; Yu et al., 2025b). Additionally, current systems discard problem-solving experiences after completion, preventing transferable knowledge accumulation across tasks. It contrasts with human problem-solving: when tackling complex tabular problems, humans methodically work through solutions, verify results, reflect on mistakes, and accumulate experiences for future strategies.
@@
To address these limitations, we propose MAPLE  (Multi-agent Adaptive Planning with Long-term mEmory), a novel framework that mimics human problem-solving through specialized cognitive agents in a feedback-driven loop. As illustrated in Figure 1, MAPLE  decomposes reasoning into distinct stages: reasoning, verification, reflection, and memory evolution, each managed by a dedicated agent. Our framework implements a feedback-driven cycle with the Solver conducting iterative reasoning, the Checker performing quality assessment, the Reflector diagnosing errors and suggesting improvements, and the Archiver managing long-term memory for cross-task learning. This architecture enables dynamic adaptation both within tasks and across similar problems, mirroring human cognitive processes.
@@ 2.1 Overview
Solver $\mathcal{T}$ or $t^{\prime},q,\tau,\mathcal{M},r$ $t^{\prime}$ or $a_{s}$ Progressive reasoning with real-time environmental feedback
[…]
@@
Inspired by human problem-solving processes, we propose MAPLE  (Multi-agent Adaptive Planning with Long-term mEmory), a novel framework that addresses a critical limitation in existing systems: their inability to adapt, reflect, and learn from experience. As illustrated in Figure 1 and formalized in Algorithm 1, MAPLE  decomposes reasoning into four distinct stages, each managed by a dedicated agent that fulfills a specialized cognitive function.
@@ 2.2.1 Solver
The Solver agent $(\mathcal{S})$ leverages the ReAct paradigm (Yao et al., 2023) to establish genuine environmental interaction with tabular data. After each table operation, the Solver reassesses the environment to determine whether to perform additional manipulations or derive an answer. This iterative reasoning-acting cycle enables strategic adaptation based on real-time feedback from the manipulated table state.
Formally, the Solver operates according to the following process, given input $(I,q,\tau,\mathcal{M},r)$:
$\lambda=\pi_{\text{solver}}(I,q,\tau,\mathcal{M},r)$ (1)
@@
A distinctive feature of our approach is the integration of dual memory systems. The memory input is defined as =$\mathcal{M}=\{\mathcal{M}_{w},\mathcal{M}_{l}\}$, where working memory $\mathcal{M}_{w}$ contains Reflector feedback ($d,p$) providing diagnostic insights when prior attempts failed. Long-term memory $\mathcal{M}_{l}$ retrieves relevant historical experiences ($\mathcal{N}_{\text{solver}}$), including similar questions, proven strategies and common pitfalls, creating a knowledge repository that enriches the reasoning process.
@@ 2.2.3 Reflector
3:$Finished\leftarrow\text{False}$triangleright Initialization flag
4:$\tau\leftarrow[]$triangleright Initialize Solver operation history
@@
6:     $\mathcal{N}_{\text{solver}}\leftarrow\pi_{\text{mem-retrieve}}^{\text{solver}}(\mathcal{T},q\mid\mathcal{M}_{l},k,\delta)$triangleright Retrieve neighbor memories, Eq. 5
7:     $t^{\prime}$ or $a_{s}\leftarrow\pi_{\text{solver}}(\mathcal{I},q,\tau,r,\mathcal{N}_{\text{solver}},d,p)$triangleright Operated table or answer of Solver, Eq. 1
8:     $\tau.\text{append(LogSolverOperation($\mathcal{M}_{w}))$}$triangleright Log current operation in operation history
[…]
10:         $\mathcal{F}\leftarrow\pi_{\text{checker}}(\mathcal{T},q,a_{s})$triangleright Evaluate answer, Eq. 2
[…]
12:              $a_{m}\leftarrow a_{s}$triangleright Accept Solver answer
@@
15:              $(d,p)\leftarrow\pi_{\text{reflector}}(\mathcal{T},q,\tau,a_{s},\mathcal{F})$triangleright Diagnose and provide fix plan, Eq. 3
16:              $\mathcal{M}_{w}\leftarrow\text{UpdateWorkingMemory}(d,p)$triangleright Inject feedback for adaptive refinement
@@
19:         Continue reasoning with updated $t^{\prime}$triangleright Continue ReAct loop
@@
21:     $r\leftarrow r-1$triangleright Decrease remaining attempts
@@
23:return $a_{m}$triangleright Output model prediction
@@ 2.2.4 Archiver
The Archiver agent $(\mathcal{A})$ introduces experiential learning through structured memory management. This agent implements knowledge retention, retrieval, and evolution mechanisms that enable continual improvement across reasoning tasks, with detailed working memory and long-term memory mechanisms described in Section B.1. The Archiver operates through three complementary modules:
@@
Through this sophisticated memory management approach, MAPLE  transcends traditional systems that restart reasoning from scratch on each task. Instead, our framework builds an evolving knowledge repository that improves performance across similar problems through continuous refinement. For a more detailed algorithm describing this memory evolution process, please see Appendix C. The complete Archiver prompt is provided in Appendix F.
Having defined the specialized roles and interactions of each agent in our framework, we now formalize the complete reasoning procedure that orchestrates their collaborative operation.  Algorithm 1 presents the adaptive reasoning loop of MAPLE, illustrating how multiple agents coordinate through a feedback-driven cycle to progressively refine reasoning strategies. This algorithm demonstrates several key innovations absent in traditional approaches: (1) iterative refinement through verification and reflection, (2) dynamic adaptation based on structured feedback, and (3) integration of experiential knowledge from similar problems.
@@ 3.1 Experimental Setup
Model LLaMA 3 Qwen 2.5 LLaMA 3 Qwen 2.5
End-to-End QA 45.58 35.93 75.49 63.49
Few-Shot QA 58.54 37.66 74.73 67.49
Chain-of-Thought 65.75 63.12 75.89 77.67
Binder (Cheng et al., 2023) 65.26 63.64 74.54 80.20
Dater (Ye et al., 2023) 62.18 61.94 81.12 78.46
CoTable (Wang et al., 2024b) 64.80 67.87 83.25 81.96
ReAcTable (Zhang et al., 2023) 63.07 63.74 81.49 82.06
MAPLE (Ours) 74.01 73.39 90.66 86.02
↑ 8.26 ↑ 5.52 ↑ 7.41 ↑ 3.96
Table 2: Table reasoning accuracy on WikiTQ and TabFact using Qwen2.5-72B and LLaMA3.3-70B. Bold denotes the best performance and underline denotes the second-best performance in each column. Red arrows indicate improvements over the strongest baseline.
@@
We evaluate our approach on two standard benchmarks: (1) WikiTableQuestions (WiKiTQ(Pasupat and Liang, 2015): A widely used benchmark dataset for studying question answering over structured tables. It contains 14,149 question-answer pairs for training and 4,344 for testing, collected from 421 Wikipedia tables. The questions require different levels of reasoning, and the answers can be single values, lists of values, or derived results that are not explicitly present in the table. (2) TabFact (Chen et al., 2020): A benchmark for fact verification over tabular data, consisting of natural language statements paired with tables from diverse domains. Each statement is labeled as either entailed ("yes") or refuted ("no") based on the table content. The test set includes 2,024 statements across 298 tables, requiring models to perform complex reasoning to verify factual accuracy.
@@
We compare our multi-agent framework against three categories of baseline approaches: (1) Standard Reasoning, where the model directly generates answers from the table and question. This includes End-to-End QA, which outputs the answer in a single step. Few-Shot QA, which adds example (Table, Question, Answer) triplets to guide the model. Chain-of-Thought (Wei et al., 2023), which encourages the model to explain its reasoning process before answering. (2) Program-Based Reasoning, which guide the model to produce executable code for answering. Binder (Cheng et al., 2023) prompts the model to generate Python or SQL code. Dater (Ye et al., 2023) breaks down the question and table into smaller parts for easier processing. (3) ReAct-Based Reasoning: This approach integrates reasoning and acting in an iterative process, using external tools to assist decision-making. Chain-of-Table (Wang et al., 2024b) dynamically constructs intermediate tables to support reasoning. ReAcTable (Zhang et al., 2023) follows this paradigm by integrating SQL and Python executions to generate intermediate results and refine reasoning steps.
@@
We conduct our experiments using two state-of-the-art LLMs: LLaMA3.3-70B-Instruct (Grattafiori et al., 2024) [^2] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct and Qwen2.5-72B-Instruct (Qwen et al., 2025) [^3] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct. All models run on two NVIDIA A100 GPUs. The tabular input is converted into markdown format before being passed to the LLMs. We use in-context prompting by including task-specific examples, which are provided in Appendix F. Default decoding parameters are used throughout. For all baseline methods, we follow their original settings to ensure optimal performance.

### v2

@@ Abstract
Table-based question answering requires complex reasoning capabilities that current LLMs struggle to achieve with single-pass inference. Existing approaches, such as Chain-of-Thought reasoning and question decomposition, lack error detection mechanisms and discard problem-solving experiences, contrasting sharply with how humans tackle such problems. In this paper, we propose MAPLE  (Multi-agent Adaptive Planning with Long-term mEmory), a novel framework that mimics human problem-solving through specialized cognitive agents working in a feedback-driven loop. MAPLE  integrates 4 key components: (1) a Solver using the ReAct paradigm for reasoning, (2) a Checker for answer verification, (3) a Reflector for error diagnosis and strategy correction, and (4) an Archiver managing long-term memory for experience reuse and evolution. Experiments on WiKiTQ and TabFact demonstrate significant improvements over existing methods, achieving state-of-the-art performance across multiple LLM backbones. [^1] Dataset and code are available at https://github.com/bettyandv/MAPLE-table-reasoning
@@ 1 Introduction
Tables represent one of the most prevalent forms of semi-structured data, organizing information systematically across domains ranging from scientific research to business analytics (10.1145/3626772.3661384). However, answering questions over tables presents unique challenges, requiring multi-step reasoning over structured data, recognition of implicit relationships between cells, and precise contextual interpretation (Lu_2025). These challenges make table-based question answering (QA) particularly difficult for Large Language Models (LLMs), as they must navigate tabular data structure while performing sophisticated reasoning to derive accurate answers, capabilities that current LLMs struggle to achieve with single-pass inference.
Existing table reasoning frameworks exhibit several limitations. Single-forward-pass methods (cheng_binding_2023; ye_large_2023) lack error detection mechanisms, allowing mistakes to propagate through solutions. ReAct-based approaches (wang_chain--table_2024; zhang_reactable_2023) provide environmental feedback but lack systematic verification. On the other hand, multi-agent approaches primarily focus on output refinement rather than comprehensive reasoning improvement (ye_large_2023; yu2025tablecriticmultiagentframeworkcollaborative). Additionally, current systems discard problem-solving experiences after completion, preventing transferable knowledge accumulation across tasks. It contrasts with human problem-solving: when tackling complex tabular problems, humans methodically work through solutions, verify results, reflect on mistakes, and accumulate experiences for future strategies.
@@
To address these limitations, we propose MAPLE  (Multi-agent Adaptive Planning with Long-term mEmory), a novel framework that mimics human problem-solving through specialized cognitive agents in a feedback-driven loop. As illustrated in Figure˜1, MAPLE  decomposes reasoning into distinct stages: reasoning, verification, reflection, and memory evolution, each managed by a dedicated agent. Our framework implements a feedback-driven cycle with the Solver conducting iterative reasoning, the Checker performing quality assessment, the Reflector diagnosing errors and suggesting improvements, and the Archiver managing long-term memory for cross-task learning. This architecture enables dynamic adaptation both within tasks and across similar problems, mirroring human cognitive processes.
@@
2 Related Work
Multi-agent System
Multi-agent systems (MAS) have gained attention for leveraging collective intelligence in complex tasks. Current applications span problem-solving domains such as software development (li2023camel; du2023improvingfactualityreasoninglanguage; qian2024chatdevcommunicativeagentssoftware; huang2024agentcodermultiagentbasedcodegeneration), embodied robotic coordination (dasgupta2023collaboratinglanguagemodelsembodied; mandi2023rocodialecticmultirobotcollaboration; zhang2024buildingcooperativeembodiedagents; yu2025conavgptmultirobotcooperativevisual), and scientific Debate (du2023improvingfactualityreasoninglanguage; Xiong_2023; tang2024medagentslargelanguagemodels). World simulation applications include social behavior modeling (park2022socialsimulacracreatingpopulated; gao2023s3socialnetworksimulationlarge; chen2025multiagentconsensusseekinglarge), policy simulation (xiao2023simulatingpublicadministrationcrisis; hua2024warpeacewaragentlarge), economic forecasting (horton2023largelanguagemodelssimulated; li2023tradinggptmultiagentlayeredmemory; zhao2024competeaiunderstandingcompetitiondynamics), and gaming (light2023avalonbenchevaluatingllmsplaying; wang2023avalonsgamethoughtsbattle; xu2024languageagentsreinforcementlearning). Several MAS frameworks, including MetaGPT (hong2024metagptmetaprogrammingmultiagent), CAMEL (li2023camel), and AutoGen (wu2023autogenenablingnextgenllm), have emerged to facilitate the implementation of these systems.
Memory Mechanisim for Agent
Memory enables agents to perform coherent, context-aware reasoning across extended tasks. In multi-agent systems, memory serves as the cognitive foundation for retaining observations, decisions, and interaction histories—critical elements for consistent collaboration and adaptation over time (sumers2024cognitivearchitectureslanguageagents).
Memory formats determine how content is stored and utilized, with several prominent approaches emerging in recent work. Natural language representations offer semantic richness and interpretability, as demonstrated in Reflexion (shinn2023reflexionlanguageagentsverbal), Voyager (wang2023voyageropenendedembodiedagent), and Generative Agents (park2023generativeagentsinteractivesimulacra). Vector embeddings enable efficient similarity-based retrieval, a technique central to systems like MemoryBank (zhong2023memorybankenhancinglargelanguage), A-MEM (xu2025amemagenticmemoryllm), and ChatDev (qian2024chatdevcommunicativeagentssoftware). Meanwhile, structured formats support symbolic reasoning and precise queries, approaches adopted by ChatDB (hu2023chatdbaugmentingllmsdatabases) and DB-GPT (zhou2023llmdba). Beyond storage formats, recent advances have introduced innovative management strategies, including complete interaction storage (zhong2023memorybankenhancinglargelanguage; modarressi2024retllmgeneralreadwritememory), cache-like designs (packer2024memgptllmsoperatingsystems), and controller-based architectures (wang2025scmenhancinglargelanguage) that dynamically prioritize and maintain relevant information during extended reasoning processes.
Table Reasoning
Research on table reasoning can be broadly classified into fine-tuning-based and prompting-based methods. Fine-tuning approaches like TAPAS (herzig_tapas_2020), Pasta (gu_pasta_2022), TUTA (wang_tuta_2021), and TAPEX (liu_tapex_2022) adapt pre-trained language models to encode table semantics through specialized training objectives. Other works focus on improving alignment between natural language queries and structured data (eisenschlos-etal-2020-understanding; jiang-etal-2022-omnitab). Despite their effectiveness, these methods typically require extensive annotated data and feature static reasoning processes without adaptive correction mechanisms.
Prompting-based methods leverage LLMs with minimal training data requirements. Techniques like Chain-of-Thought (wei2023chainofthoughtpromptingelicitsreasoning), Least-to-Most (zhou2023leasttomostpromptingenablescomplex) and Dater (ye_large_2023) perform reasoning by decomposing tasks into explicit steps. Table-specific adaptations include Binder (cheng_binding_2023), Chain-of-Table (wang_chain--table_2024), ReAcTable (zhang_reactable_2023), and Table-Critic (yu2025tablecriticmultiagentframeworkcollaborative), incorporating agent collaboration or ReAct-style reasoning.
Existing table reasoning methods exhibit significant limitations across key design dimensions. While some implement multi-agent architectures or ReAct-based reasoning, none integrates all critical components: dynamic planning, reflection mechanisms, self-refinement, and long-term memory. A detailed comparison of these methods is provided in Appendix˜A. Our proposed MAPLE  framework addresses these gaps by combining collaborative verification, adaptive planning, and evolving memory structures to achieve more robust and accurate table reasoning.
@@ 3.1 Overview
Solver $\mathcal{T}$ or $^{\prime},q,\tau,r,\mathcal{M} tt^{\prime}$ or $a_{s}$ Progressive reasoning with real-time environmental feedback
Checker $\mathcal{T},q,a_{s}$ $\mathcal{F}$ Verifies answer type, format and evidence grounding
@@
Inspired by human problem-solving processes, we propose MAPLE  (Multi-agent Adaptive Planning with Long-term mEmory), a novel framework that addresses a critical limitation in existing systems: their inability to adapt, reflect, and learn from experience. As illustrated in Figure˜1 and formalized in Algorithm˜1, MAPLE  decomposes reasoning into four distinct stages, each managed by a dedicated agent that fulfills a specialized cognitive function.
@@ 3.2.1 Solver
The Solver agent $(\mathcal{S})$ leverages the ReAct paradigm (yao2023reactsynergizingreasoningacting) to establish genuine environmental interaction with tabular data. After each table operation, the Solver reassesses the environment to determine whether to perform additional manipulations or derive an answer. This iterative reasoning-acting cycle enables strategic adaptation based on real-time feedback from the manipulated table state.
Formally, the Solver operates according to the following process, given input $(I,q,\tau,r,\mathcal{M})$:
$\lambda=\pi_{\text{solver}}(I,q,\tau,r,\mathcal{M})$ (1)
@@
A distinctive feature of our approach is the integration of dual memory systems. The memory input is defined as $\mathcal{M}=\{\mathcal{M}_{w},\mathcal{M}_{l}\}$, where working memory $\mathcal{M}_{w}$ contains Reflector feedback ($d,p$) providing diagnostic insights when prior attempts failed. Long-term memory $\mathcal{M}_{l}$ retrieves relevant historical experiences ($\mathcal{N}_{\text{solver}}$), including similar questions, proven strategies and common pitfalls, creating a knowledge repository that enriches the reasoning process.
@@ 3.2.3 Reflector
[…]
@@
6:  $\mathcal{N}_{\text{solver}}\leftarrow\pi_{\text{mem-retrieve}}^{\text{solver}}(\mathcal{T},q\mid\mathcal{M}_{l},k,\delta)$ $\triangleright$ Retrieve neighbor memories, Eq. 5
7:  $t^{\prime}$ or $a_{s}\leftarrow\pi_{\text{solver}}(I,q,\tau,r,\mathcal{M})$ $\triangleright$ Operated table or answer of Solver, Eq. 1
8:  $\tau.\text{append(LogSolverOperation($\mathcal{M}_{w}))$}$ $\triangleright$ Log current operation in operation history
9:  if $a_{s}\neq\texttt{<NOT\_READY>}$ then
10:   $\mathcal{F}\leftarrow\pi_{\text{checker}}(\mathcal{T},q,a_{s})$ $\triangleright$ Evaluate answer, Eq. 2
11:   if $\mathcal{F}.\text{total\_score}==\texttt{FULL\_SCORE}$ then
12:     $a_{m}\leftarrow a_{s}$ $\triangleright$ Accept Solver answer
@@
15:     $(d,p)\leftarrow\pi_{\text{reflector}}(\mathcal{T},q,\tau,a_{s},\mathcal{F})$ $\triangleright$ Diagnose and provide fix plan, Eq. 3
16:     $\mathcal{M}_{w}\leftarrow\text{UpdateWorkingMemory}(d,p)$ $\triangleright$ Inject feedback for adaptive refinement
@@
19:   Continue ReAct reasoning on updated table $t^{\prime}$ $\triangleright$ Continue ReAct loop
@@
21:  $r\leftarrow r-1$ $\triangleright$ Decrease remaining attempts
@@
23:return $a_{m}$ $\triangleright$ Output model prediction
@@ 3.2.4 Archiver
The Archiver agent $(\mathcal{A})$ introduces experiential learning through structured memory management. This agent implements knowledge retention, retrieval, and evolution mechanisms that enable continual improvement across reasoning tasks, with detailed working memory and long-term memory mechanisms described in Section˜B.1. The Archiver operates through three complementary modules:
@@
Through this sophisticated memory management approach, MAPLE  transcends traditional systems that restart reasoning from scratch on each task. Instead, our framework builds an evolving knowledge repository that improves performance across similar problems through continuous refinement. For a more detailed algorithm describing this memory evolution process, please see Appendix˜C. The complete Archiver prompt is provided in Appendix F.
Having defined the specialized roles and interactions of each agent in our framework, we now formalize the complete reasoning procedure that orchestrates their collaborative operation.  Algorithm˜1 presents the adaptive reasoning loop of MAPLE, illustrating how multiple agents coordinate through a feedback-driven cycle to progressively refine reasoning strategies. This algorithm demonstrates several key innovations absent in traditional approaches: (1) iterative refinement through verification and reflection, (2) dynamic adaptation based on structured feedback, and (3) integration of experiential knowledge from similar problems.
@@ 4.1 Experimental Setup
Model GPT-4o-mini LLaMA3.3-70B Qwen2.5-72B Average GPT-4o-mini LLaMA3.3-70B Qwen2.5-72B Average
End-to-End QA 46.64 45.58 35.93 42.72 71.69 75.49 63.49 70.22
Few-Shot QA 57.16 58.54 37.66 51.12 70.16 74.73 67.49 70.79
Chain-of-Thought 62.96 65.75 63.12 63.94 71.94 75.89 77.67 75.17
Binder (cheng_binding_2023) 60.24 65.26 63.64 63.05 75.45 74.54 80.20 76.73
Dater (ye_large_2023) 61.13 62.18 61.94 61.75 75.39 81.12 78.46 78.32
Chain-of-Table (wang_chain--table_2024) 62.98 64.80 67.87 65.22 79.05 83.25 81.96 81.42
ReAcTable (zhang_reactable_2023) 62.79 63.07 63.74 63.20 79.18 81.49 82.06 80.91
MAPLE (Ours) 67.13 74.01 73.39 71.51 81.62 90.66 86.02 86.10
↑ 4.15 ↑ 8.26 ↑ 5.52 ↑ 6.29 ↑ 2.44 ↑ 7.41 ↑ 3.96 ↑ 4.68
Table 2: Table reasoning accuracy on WikiTQ and TabFact using Qwen2.5-72B, LLaMA3.3-70B, and GPT-4o-mini. Bold denotes the best performance and underline denotes the second-best performance in each column. Red arrows indicate improvements over the strongest baseline.
@@
We evaluate our approach on two standard benchmarks: (1) WikiTableQuestions (WiKiTQ(pasupat_compositional_2015): A widely used benchmark dataset for studying question answering over structured tables. It contains 14,149 question-answer pairs for training and 4,344 for testing, collected from 421 Wikipedia tables. The questions require different levels of reasoning, and the answers can be single values, lists of values, or derived results that are not explicitly present in the table. (2) TabFact (chen2020tabfactlargescaledatasettablebased): A benchmark for fact verification over tabular data, consisting of natural language statements paired with tables from diverse domains. Each statement is labeled as either entailed ("yes") or refuted ("no") based on the table content. The test set includes 2,024 statements across 298 tables, requiring models to perform complex reasoning to verify factual accuracy.
@@
We compare our multi-agent framework against three categories of baseline approaches: (1) Standard Reasoning, where the model directly generates answers from the table and question. This includes End-to-End QA, which outputs the answer in a single step. Few-Shot QA, which adds example (Table, Question, Answer) triplets to guide the model. Chain-of-Thought (wei2023chainofthoughtpromptingelicitsreasoning), which encourages the model to explain its reasoning process before answering. (2) Program-Based Reasoning, which guide the model to produce executable code for answering. Binder (cheng_binding_2023) prompts the model to generate Python or SQL code. Dater (ye_large_2023) breaks down the question and table into smaller parts for easier processing. (3) ReAct-Based Reasoning: This approach integrates reasoning and acting in an iterative process, using external tools to assist decision-making. Chain-of-Table (wang_chain--table_2024) dynamically constructs intermediate tables to support reasoning. ReAcTable (zhang_reactable_2023) follows this paradigm by integrating SQL and Python executions to generate intermediate results and refine reasoning steps.
@@
We conduct our experiments using 3 state-of-the-art LLMs: GPT-4o-mini [^2] https://platform.openai.com/docs/models/gpt-4o-mini, LLaMA3.3-70B-Instruct (grattafiori2024llama3herdmodels) [^3] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct and Qwen2.5-72B-Instruct (qwen2025qwen25technicalreport) [^4] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct. All models run on two NVIDIA A100 GPUs. The tabular input is converted into markdown format before being passed to the LLMs. We use in-context prompting by including task-specific examples, which are provided in Appendix˜F. Default decoding parameters are used throughout. For all baseline methods, we follow their original settings to ensure optimal performance.